# VT-Former: A Transformer-based Vehicle Trajectory Prediction Approach For Intelligent Highway Transportation Systems

### v2

@@ Abstract
Enhancing roadway safety and traffic management has become an essential focus area for a broad range of modern cyber-physical systems and intelligent transportation systems. Vehicle Trajectory Prediction is a pivotal element within numerous applications for highway and road safety. These applications encompass a wide range of use cases, spanning from traffic management and accident prevention to enhancing work-zone safety and optimizing energy conservation. The ability to implement intelligent management in this context has been greatly advanced by the developments in the field of Artificial Intelligence (AI), alongside the increasing deployment of surveillance cameras across road networks. In this paper, we introduce a novel transformer-based approach for vehicle trajectory prediction for highway safety and surveillance, denoted as VT-Former. In addition to utilizing transformers to capture long-range temporal patterns, a new Graph Attentive Tokenization (GAT) module has been proposed to capture intricate social interactions among vehicles. Combining these two core components culminates in a precise approach for vehicle trajectory prediction. Our study on three benchmark datasets with three different viewpoints demonstrates the State-of-The-Art (SoTA) performance of VT-Former in vehicle trajectory prediction and its generalizability and robustness. We also evaluate VT-Former’s efficiency on embedded boards and explore its potential for vehicle anomaly detection as a sample application, showcasing its broad applicability.
Index Terms:
Trajectory Prediction, Road Safety, Intelligent Transportation, Transformer, Cyber-physical Systems, Artificial Intelligence
I Introduction
@@
In the realm of Cyber-Physical Systems (CPS) and its intersection with Computer Vision, the subject of Vehicle Trajectory Prediction (VTP) has emerged as a point of significance. VTP is utilized within various higher-level applications among which are traffic management and control [27, 52, 36, 39], accident and collision avoidance [17, 25, 40], workzone safety [27, 62, 64], route planning [50, 65], and vehicle anomaly detection [48, 26, 19, 24], etc.
@@
The rapid advancements in AI and the widespread availability of surveillance cameras have opened up promising avenues and introduced novel techniques for advancing vehicle trajectory prediction methods [13, 20, 32]. VTP is defined as a process that utilizes the historical vehicle movement to forecast the future path, position, and behavior of a vehicle over a specific time horizon. While many studies focus on trajectory prediction from the perspective of autonomous driving and autonomous vehicles [57, 18, 6], some studies predominantly approach it from surveillance angles, including the bird’s eye view, high-angle, and eye-level perspectives [4, 27]. In this paper, we focus on the surveillance angles in the context of highway safety.
There are several challenges and requirements associated with VTP, among which three stand out. Firstly, there are real-time constraints for VTP. Many applications of trajectory prediction, such as traffic management or vehicle anomaly detection, require real-time predictions, which add complexity in terms of computational resources and speed. Balancing the trade-off between model accuracy and computational efficiency is a challenge [43, 33, 4]. Secondly, vehicles interact with each other and with pedestrians, cyclists, and other road users. Predicting how these interactions will unfold is complex, especially when combined with the uncertain dynamics of these moving objects. Another substantial challenge lies in the limited flexibility of most proposed methods, as they often lack the adaptability to generalize effectively across diverse setups primarily tailored for specific dataset conditions. Consequently, the inherent variability of environmental factors, such as camera angles and lighting conditions, can significantly impair their performance and overall robustness.
@@
In this paper, we present VT-Former, an innovative approach that leverages the power of transformers for vehicle trajectory prediction in roadways. VT-Former also introduces Graph Attentive Tokenization (GAT) to capture the underlying social interactions among moving vehicles within the scene. It allows us to take full advantage of spatial and temporal features, resulting in a holistic and accurate approach to modeling the complex dynamics of vehicle movements in diverse and dynamic environments. Integrating these key components, GAT, and the decoder-only transformer, within VT-Former, offers a State-of-the-Art (AotA) algorithmic approach for vehicle trajectory prediction.
To demonstrate the benefits of VT-Former, we perform extensive experimentation using three CPS-relevant benchmark datasets: NGSIM [11, 12], CHD High-angle and CHD Eye-level [28] datasets, to comprehensively evaluate the performance of the VT-Former model from a multidimensional perspective. The evaluation encompassed three distinct angle variations corresponding to each dataset. VT-Former demonstrated exceptional performance, surpassing the efficacy of prior models. Notably, on the NGSIM dataset, our design achieved an Average Displacement Error (ADE) of 0.82 meters and a Final Displacement Error (FDE) of 1.86 meters, establishing a new state-of-the-art (SoTA) standard for the prediction of vehicle trajectories. Additionally, VT-Former demonstrated competitive results on both the CHD Eye-level and High-angle datasets.
To illustrate the adaptability of VT-Former for CPS applications, we assessed its performance on an embedded system. We delved into a thorough analysis, comparing the effects of using shortened historical trajectory data against extended historical horizons. This analysis revealed the model’s precision in forecasting future trajectories. Such an attribute is essential across numerous applications, especially in contexts like vehicle anomaly detection where prompt responses are paramount. Emphasizing VT-Former’s value in a vital CPS application aimed at boosting highway safety, we introduce an intuitive yet potent driving anomaly detection method. Drawing inspiration from predictive-based anomaly detection techniques in domains such as cyber security and public safety, we utilized VT-Former’s state-of-the-art accuracy to gauge its proficiency in predictive-based driving anomaly detection. We validated our approach using a dataset of anomalous driving behaviors, generated through an adversarial anomalous trajectory creation algorithm from NGSIM, highlighting its promising capability to tackle real-world issues.
@@
  • Introducing VT-Former. A novel approach for vehicle trajectory prediction for highway safety applications.
@@
  • Exploring the influence of reduced observation horizons on model accuracy, underscoring its effectiveness in scenarios demanding rapid trajectory prediction.
  • Assessing the real-time efficiency of the model by executing it on an embedded platform.
  • Demonstrating the practical utility of VT-Former by implementing a straightforward driving anomaly detection algorithm and evaluating it on adversarially generated anomaly scenarios.
II Related Works
@@
In recent years, intelligent transportation systems (ITS) have emerged as a focal point of extensive research endeavors within the scientific community [1, 53, 16, 44, 58]. Modern edge-based Cyber-Physical Systems are increasingly integrating deep learning-based prediction mechanisms [47, 4, 38]. For many CPS applications, path prediction is crucial as it helps in foreseeing future actions and locations beneficial for many downstream tasks. These predictive tools operate in real-time at local edge nodes [21, 27]. Such methodologies can be utilized for forecasting traffic densities [22, 9], improving the safety of work zones [54], and promoting autonomous vehicle technologies [55]. These predictive models also play a critical role in enabling advancements in smart monitoring and surveillance for enhanced security [49], improving pedestrian and worker safety in urban environments [42, 61], facilitating the identification and forecasting of anomalies within complex systems [41], and optimizing the design of transportation routes for increased efficiency and resource allocation [7].
@@
Traditional VTP methods, such as those based on Long Short-Term Memory (LSTM) networks, have been explored by Alahi et al. [2]. While effective, these methods often lack the flexibility and scalability of newer architectures. Graph Neural Networks (GNNs) have emerged as a powerful tool for trajectory prediction [35, 15], capitalizing on their ability to model relational data and complex interactions between entities. Specifically, Graph Convolutional Networks (GCN) [10, 55] allows for localized feature learning on graphs, making them apt for capturing spatial relationships between vehicles. The Graph Isomorphism Network (GIN) [42] further extends this capability by distinguishing different graph structures, enabling more precise trajectory predictions by considering nuanced vehicular interactions [66, 60, 31].
@@
In recent studies, there has been a notable shift from Long Short-Term Memory (LSTM) and Convolutional Neural Networks (CNNs) to Transformers, as introduced by Vaswani et al. [59]. Transformers employ self-attention mechanisms for the analysis of sequential data making them suitable for many applications such as computer vision [30], natural language processing [56], molecular chemistry [51], etc. Simultaneously, there have been endeavors to harness the power of Transformers in predicting trajectories within the realm of autonomous vehicles, enabling a better understanding of the temporal evolution of object movements [67, 23], yet these methods have yet to be applied in the context of highway surveillance and safety applications.
@@
Despite the multitude of trajectory prediction algorithms proposed over the past decade, a limited number are explicitly designed for highway environments, with even fewer tailored for resource-constrained edge devices [5, 27]. Furthermore, the field of highway trajectory prediction is hampered by the scarcity of comprehensive datasets. Some recent works such as [28] have attempted to mitigate this and fill the gap. However, there remains a distinct requirement for specialized datasets tailored to highway-based applications. In general, there is a noticeable disparity between research endeavors and real-world applications in the realm of trajectory prediction based on already installed surveillance cameras. This underscores the potential for research and further development in this domain.
III VT-Former Formulation and Methodology
@@
At its core, VT-Former is a Transformer based approach for vehicle pass prediction in roadways. Transformers are renowned for their exceptional generalization capabilities [59], enabling them to discern intricate long-term patterns within data and excel in tasks featuring extensive and diverse datasets. While the original Transformer model was computationally demanding and exhibited slower performance as it had both encoder and decoder components, numerous optimizations and model adaptations have been developed to enhance real-time processing [29, 8] by choosing only encoder or decoder as the core of their models. This has facilitated the utilization of transformers in applications necessitating low-latency responses. These attributes motivated our choice of the decoder-only transformer [8] as the core component of our algorithm.
@@
Figure 2: VT-Former at a glance: our approach begins with the application of Graph Attentive Tokenization (GAT) to create enriched token sequences that encapsulate social interactions. Subsequently, these enhanced sequences are fed through the Transformer Prediction module, which consists of 8 layers of transformer decoders, each equipped with 4 heads, generating predicted trajectories. $T_{OH}$, $TS_{i}$, $T_{PH}$, and $PC_{i}$ denote the observation horizon time, token sequence of the $i^{th}$ vehicle, prediction horizon time, and the predicted trajectory, respectively.
III-A Graph Attention Tokenization
Graph Attention Tokenization (GAT) illustrated in Fig. 1 aims to extract enriched tokens that encapsulate the social interactions between the vehicles available in the scene. In order to accentuate the relative movements of the vehicles through the observation window we define the relative trajectory with respect to the first observed location as:
@@ III-A Graph Attention Tokenization
$C_{i}$ and $\Delta C_{i}$ are concatenated to form the input of the fully connected layer which expands the dimension of the input features by a factor of $2$. In the next step, we leverage Graph Isomorphism Network (GIN) [66] to capture the relationships between subjects available in the scene. GINs have proven highly effective in the context of processing graph-structured data, where nodes and edges signify entities and their relationships, respectively. In our formulation, each vehicle will be considered as a node in a fully connected graph. We chose to have a fully connected graph so that the Graph Network does not have any predefined biases and all relationships are learned by training. Another benefit of having a fully connected graph is that by performing just one step graph operation, the information will be propagated over all nodes and further steps are not necessary which makes the design more efficient for CPS applications. Inspired by [4], we design separate aggregation networks for the node feature and the neighbor features. The node feature aggregator is an MLP network with one hidden layer:
@@
III-B Transformer Predictor Module
@@
where $W_{3}$, $b_{3}$, $W_{4}$, and $b_{4}$ are the parameters of the two fully connected layers. The Transformer Predictor Module consists of 8 decoder layers stacked on top of each other. We chose the dimension of the Feed Forward Network (FFN) to be a size of 256. Within our design, as detailed in Sec. III-A, we incorporated inter-vehicle attention in the tokenization process, although temporal attention among tokens was not considered. The multi-head attention mechanism in transformers is an ensemble of distinct attention mechanisms, with each head specializing in particular patterns. This adaptability empowers transformers to effectively capture a wide spectrum of data dependencies in the temporal dimension, ultimately improving the accuracy of predicted trajectories. We choose an autoregressive strategy where the model generates the sequence of predicted locations one step at a time in an ordered manner, where each prediction is influenced by the previous predictions leading to more accurate prediction. To prevent the Transformer Predictor from peeking into future information, we employ future masking while training. The masking procedure is particularly important to ensure the causality of the system in parallel training mode.
IV Experimental Setup
@@
IV-A Datasets
IV-A1 NGSIM
@@
IV-A2 Carolina’s Highway Dataset
@@
IV-B Metrics
IV-B1 Average Displacement Error (ADE)
is a metric commonly used in the field of trajectory forecasting and motion prediction. It quantifies the accuracy of predicted trajectories for moving objects, such as vehicles or pedestrians by measuring the average Euclidean distance between the ground truth coordinates ($x^{t},y^{t}$) and the predicted coordinates ($x^{\prime t},y^{\prime t}$) across a defined prediction horizon ($T_{PH}$) considering all available $N$ vehicles in the scene:

### v4

# VT-Former: An Exploratory Study on Vehicle Trajectory Prediction for Highway Surveillance through Graph Isomorphism and Transformer

@@ Abstract
Enhancing roadway safety has become an essential computer vision focus area for Intelligent Transportation Systems (ITS). As a part of ITS, Vehicle Trajectory Prediction (VTP) aims to forecast a vehicle’s future positions based on its past and current movements. VTP is a pivotal element for road safety, aiding in applications such as traffic management, accident prevention, work-zone safety, and energy optimization. While most works in this field focus on autonomous driving, with the growing number of surveillance cameras, another sub-field emerges for surveillance VTP with its own set of challenges. In this paper, we introduce VT-Former, a novel transformer-based VTP approach for highway safety and surveillance. In addition to utilizing transformers to capture long-range temporal patterns, a new Graph Attentive Tokenization (GAT) module has been proposed to capture intricate social interactions among vehicles. This study seeks to explore both the advantages and the limitations inherent in combining transformer architecture with graphs for VTP. Our investigation, conducted across three benchmark datasets from diverse surveillance viewpoints, showcases the State-of-the-Art (SotA) or comparable performance of VT-Former in predicting vehicle trajectories. This study underscores the potential of VT-Former and its architecture, opening new avenues for future research and exploration.
1 Introduction
@@
Intelligent Transportation Systems (ITS) have emerged as a significant area of interest in research, with a notable classification distinguishing between autonomous driving [15, 88, 73, 72, 22, 42, 24, 62, 19, 20, 47] and surveillance applications [61, 40, 65, 90, 81, 18]. The subject of Vehicle Trajectory Prediction (VTP), as a subset of ITS, has garnered significant attention in both areas. VTP is utilized within various higher-level applications among which are traffic management and control [36, 68, 48, 52, 80], accident and collision avoidance [25, 34, 54], workzone safety [36, 83, 86], route planning [67, 87], and vehicle anomaly detection [64, 35, 27, 33], etc.
Figure 1: A conceptual overview of the VT-Former. The input sequence, characterized by an observation horizon denoted as $T_{OH}$, undergoes tokenization via a novel graph-based tokenizer. The resultant tokens are subsequently processed through a transformer-based predictive model to forecast future trajectories.
@@
The rapid advancements in computer vision and the widespread availability of surveillance cameras have opened up promising avenues and introduced novel techniques for advancing Surveillance Vehicle Trajectory Prediction (SVTP) methods [14, 28, 43]. This study focuses on SVTP in which the historical vehicle movement is used to forecast the future path, position, and behavior of a vehicle over a specific time horizon. While many studies focus on trajectory prediction from the perspective of autonomous driving and autonomous vehicles [76, 26, 5, 56], SVTP models approach it from surveillance angles, including the bird’s eye view, high-angle, and eye-level perspectives [3, 36]. Another notable distinction between the two domains is that models in autonomous driving often predict multiple trajectories and select the most suitable one for evaluation [58, 14]. In contrast, SVTP methods typically rely on deterministic models that forecast a singular trajectory [3, 36, 46], which is more aligned with the practical requirements of highway safety scenarios.
There are several challenges and requirements associated with SVTP. Many applications such as traffic management or vehicle anomaly detection, require a balance between efficiency and accuracy [58, 44, 3]. This balance is pivotal as it directly impacts the real-world applicability and effectiveness of SVTP models in dynamic and often unpredictable environments. Another challenge is that vehicles interact with each other. Predicting how these interactions will unfold is complex, especially when combined with the uncertain dynamics of vehicle movements. Furthermore, a significant challenge stems from the design of most SVTP approaches for a bird’s-eye view, whereas highway surveillance typically employs high-angle view cameras instead of bird’s-eye views. Therefore, it is crucial to assess such models from the perspective of these more common real-world viewpoints as well.
@@
In this paper, we present VT-Former, an innovative approach that leverages the power of transformers for vehicle trajectory prediction for highway surveillance applications. VT-Former also introduces Graph Attentive Tokenization (GAT) to capture the underlying social interactions among moving vehicles within the scene. The use of graphs has demonstrated significant success in pedestrian trajectory prediction, as evidenced by prior studies [57, 3]. This exploratory study aims to assess their effectiveness in the vehicle surveillance domain. Integrating these key components, GAT, and the decoder-only transformer, within VT-Former, offers a novel approach for SVTP. Fig. 1 shows a conceptual overview of VT-Former.
To demonstrate the benefits and limitations of VT-Former, we perform extensive experimentation using three major datasets: NGSIM [12, 13], CHD High-angle and CHD Eye-level [37] datasets, to comprehensively evaluate the performance of the VT-Former model from diverse perspectives. The evaluation encompassed three distinct angle variations corresponding to each dataset. Please note that VT-Former is designed for highway surveillance, not autonomous driving, making datasets like Argoverse [84, 9], nuScenes [8], PUP [30], and Waymo [74, 21] unrelated to this work. The outcomes underscore the advantages and potential, as well as the shortcomings of employing GAT alongside transformers for enhancing trajectory prediction accuracy and efficiency. On top of comparison with SotA models, we delved into a thorough analysis, comparing the effects of shortened historical trajectory data against extended historical horizons. This analysis explores the benefits and shortcomings of using a shortened observation horizon in cases where quick response time or lower prediction latency is critical. Moreover, our ablation study demonstrates the advantages of utilizing a higher frame rate over the conventional setup, while preserving the existing ratio between the prediction horizon and the observed trajectory.
@@
  • Introducing VT-Former, exploring the combination of graph technology and transformer architecture as a new avenue and a novel approach for vehicle trajectory prediction for highway safety applications.
@@
  • Exploring the influence of reduced observation horizons as well as higher frame rate on model accuracy, underscoring its effectiveness in scenarios demanding rapid trajectory prediction.
2 Related Works
@@
In recent years, ITS have emerged as a focal point of extensive research endeavors within the scientific community [1, 69, 23, 59, 77]. Modern systems are increasingly integrating deep learning-based prediction mechanisms [63, 3, 50]. For many applications, path prediction is crucial as it helps in foreseeing future actions and locations beneficial for many downstream tasks. These predictive tools normally aim to operate in real-time [29, 36]. Such methodologies can be utilized for forecasting traffic densities [31, 10], improving the safety of work zones [70], and promoting autonomous vehicle technologies [71]. These predictive models also play a critical role in enabling advancements in smart monitoring and surveillance for enhanced security [66], improving pedestrian and worker safety in urban environments [57, 82], facilitating the identification and forecasting of anomalies within complex systems [55], and optimizing the design of transportation routes for increased efficiency and resource allocation [6].
Figure 2: VT-Former at a glance: our approach begins with the application of Graph Attentive Tokenization (GAT) to create enriched token sequences that encapsulate social interactions. Subsequently, these enhanced sequences are fed through the Transformer Prediction module, which consists of 8 layers of transformer decoders, each equipped with 4 heads, generating predicted trajectories. VT-Former is an autoregressive sequence-to-sequence model that generates the output in multiple steps. However, for easier visualization, the output tokens are shown in the last step where they are completely generated. $T_{OH}$, $TS_{i}$, $T_{PH}$, and $PC_{i}$ denote the observation horizon time, token sequence of the $i^{th}$ vehicle, prediction horizon time, and the predicted trajectory, respectively.
@@
Traditional methods, such as those based on Long Short-Term Memory (LSTM) networks, have been explored by Alahi et al. [2]. While effective, these methods often lack the flexibility and scalability of newer architectures. Graph Neural Networks (GNNs) have emerged as a powerful tool for trajectory prediction [45, 17], capitalizing on their ability to model relational data and complex interactions between entities. Specifically, Graph Convolutional Networks (GCN) [11, 71] allows for localized feature learning on graphs, making them apt for capturing spatial relationships between vehicles [53, 51]. The Graph Isomorphism Network (GIN) [57] further extends this capability by distinguishing different graph structures, enabling more precise trajectory predictions by considering nuanced vehicular interactions [89, 79, 41].
@@
In recent studies, there has been a notable shift from Long Short-Term Memory (LSTM) and Convolutional Neural Networks (CNNs) to Transformers, as introduced by Vaswani et al. [78]. Transformers employ self-attention mechanisms for the analysis of sequential data making them suitable for many applications such as computer vision [39] and natural language processing [75]. Simultaneously, there have been endeavors to harness the power of Transformers in predicting trajectories within the realm of autonomous vehicles, enabling a better understanding of the temporal evolution of object movements [91, 32]. However, transformer-based methods remain largely unexplored within highway surveillance and safety applications. Moreover, their potential synergy with graphs to more effectively capture social nuances has yet to be investigated.
@@
Despite the multitude of trajectory prediction algorithms proposed over the past decade, a limited number are explicitly designed for highway environments [4, 36]. Furthermore, the field of highway trajectory prediction is hampered by the scarcity of comprehensive datasets. Some recent works such as [37] have attempted to mitigate this and fill the gap. However, there remains a distinct requirement for specialized datasets tailored to highway-based applications. In general, there is a noticeable disparity between research endeavors and real-world applications in trajectory prediction especially when using already installed surveillance cameras. This underscores the potential for research and further development in this domain.
3 VT-Former Formulation and Methodology
@@
At its core, VT-Former is a transformer-based approach for SVTP in roadways. Transformers are renowned for their exceptional generalization capabilities [78], enabling them to discern intricate long-term patterns within data and excel in tasks featuring extensive and diverse datasets. While the original Transformer model was computationally demanding and exhibited slower performance as it had both encoder and decoder components, numerous optimizations and model adaptations have been developed to enhance their efficiency [38, 7] by choosing only encoder or decoder as the core of their models. This has facilitated the utilization of transformers in applications necessitating low-latency responses. These attributes motivated our choice of the decoder-only transformer [7] as the core component of our algorithm.
@@
3.1 Graph Attention Tokenization
Graph Attention Tokenization (GAT) illustrated in Fig. 3 aims to extract enriched tokens that encapsulate the social interactions between the vehicles available in the scene. In order to accentuate the relative movements of the vehicles through the observation window we define the relative trajectory with respect to the first observed location:
@@ 3.1 Graph Attention Tokenization
$C_{i}$ and $\Delta C_{i}$ are concatenated to form the input of the fully connected layer which expands the dimension of the input features by a factor of $2$. In the next step, we leverage Graph Isomorphism Network (GIN) [89] to capture the relationships between subjects available in the scene. GINs have proven highly effective in the context of processing graph-structured data, where nodes and edges signify entities and their relationships, respectively. In our formulation, each vehicle will be considered as a node in a fully connected graph. We chose to have a fully connected graph so that the Graph Network does not have any predefined biases and all relationships are learned by training. Another benefit of having a fully connected graph is that by performing just one step graph operation, the information will be propagated over all nodes and further steps are not necessary leading to a more efficient design. Inspired by [3], we design separate aggregation networks for the node feature and the neighbor features. The node feature aggregator is an MLP network with one hidden layer:
@@
3.2 Transformer Predictor Module
@@
where $W_{3}$, $b_{3}$, $W_{4}$, and $b_{4}$ are the parameters of the two fully connected layers.
The Transformer Predictor Module consists of 8 decoder layers stacked on top of each other. We chose the dimension of the FFN to be a size of 256. Within our design, as detailed in Sec. 3.1, we incorporated inter-vehicle attention in the tokenization process, although temporal attention among tokens was not considered. The multi-head attention mechanism in transformers is an ensemble of distinct attention mechanisms, with each head specializing in particular patterns. This adaptability empowers transformers to effectively capture a wide spectrum of data dependencies in the temporal dimension, ultimately improving the accuracy of predicted trajectories. We choose an autoregressive strategy where the model generates the sequence of predicted locations one step at a time in an ordered manner, where each prediction is influenced by the previous predictions leading to more accurate prediction. To prevent the Transformer Predictor from peeking into future information, we employ future masking while training. The masking procedure is particularly important to ensure the causality of the system in parallel training mode.
4 Experimental Setup
@@
4.1 Datasets
This study employs three datasets, each reflecting a distinct surveillance viewpoint (bird’s-eye, high-angle, and eye-level), to evaluate the performance and robustness of the VT-Former model.
4.1.1 NGSIM
@@
4.1.2 Carolina’s Highway Dataset
@@
4.2 Metrics
In alignment with previous studies, three established metrics, namely Average Displacement Error (ADE), Final Displacement Error (FDE), and Root Mean Squared Error (RMSE), are utilized to assess VT-Former’s performance.
4.2.1 Average Displacement Error (ADE)
ADE is a metric commonly used in the field of trajectory forecasting and motion prediction. It quantifies the accuracy of predicted trajectories for moving objects, such as vehicles or pedestrians by measuring the average Euclidean distance between the ground truth coordinates ($x^{t},y^{t}$) and the predicted coordinates ($x^{\prime t},y^{\prime t}$) across a defined prediction horizon ($T_{PH}$) considering all available $N$ vehicles in the scene: